Evaluate the following functions in differential operator form.

  1. D(x2+2x-3) Ans. (2x+2)
  2. D2(xe3x-e4x) Ans. 9xe3x+6e3x-16e4x
1
Expert's answer
2022-01-19T13:33:56-0500

"1.\ D\left(x^{2}+2 x+3\right)"

Since "D=\frac{d}{d x}"

So, we have,

 "\begin{aligned}

&\frac{d}{d x}\left(x^{2}+2 x+3\right) \\

&=\frac{d}{d x}\left(x^{2}\right)+\frac{d}{d x}(2 x)+\frac{d}{d x}(3) \\

&=2 x+2 \\

&2.\ D^{2}\left(x e^{3 x}-e^{4 x}\right)

\end{aligned}"  

Since "D=\frac{d}{d x}"

So, we have,

 "\begin{aligned}

&\frac{d^{2}}{d x^{2}}\left(x e^{3 x}-e^{4 x}\right) \\

&\frac{d}{d x}\left(\frac{d}{d x}\left(x e^{3 x}-e^{4 x}\right)\right)

\end{aligned}"

"=\frac{d}{d x}\left(\frac{d}{d x} x e^{3 x}-\frac{d}{d x} e^{4 x}\right)"  

Recall the product rule of differentiation

 "U V=V d u+U d v"

By comparison let "u=x, v=e^{3 x}"

Upon differentiation. We have,

 "\begin{aligned}

&\frac{d}{d x}\left(e^{3 x}+3 x e^{3 x}-4 e^{4 x}\right) \\

&=\frac{d}{d x}\left(e^{3 x}\right)+3 \frac{d}{d x}\left(x e^{3 x}\right)-4 \frac{d}{d x}\left(e^{4 x}\right) \\

&=3 e^{3 x}+3\left(e^{3 x}+3 x e^{3 x}\right)-4\left(4 e^{4 x}\right) \\

&=3 e^{3 x}+3 e^{3 x}+9 x e^{3 x}-16 e^{4 x} \\

&=9 x e^{3 x}+6 e^{3 x}-16 e^{4 x}

\end{aligned}"